Three times parents age of a father is equal to eight time age of his son. Eight years hence the father will twice as old as his son at that time. what are the paresant ages???
Answer by stanbon(75887)

Three times parents age of a father is equal to eight time age of his son. Eight years hence the father will twice as old as his son at that time. what are the paresant ages???
------------------------
Equation:
3f = 8s
f+8 = 2(s+8)
-----
Solve for "s" and substitute to solve for "f":
s = (3/8)f
------
Sustituting you get:
f + 8 = 2((3/8)f+8)
f + 8 = (1/2)f + 16
-----
(1/2)f = 8
f = 16 yrs
Note: That age for the father doesn't make much sense.
-----
Cheers,
Stan H.